an object of mass 90 kg is being moved over an inclined plane of two angles 30 and 60. Calculate the force required in each case to achieve the acceleration of 1 m/s^2
1
Expert's answer
2019-08-09T10:49:27-0400


Newton’s second law:


"\vec{N}+m\vec{g}+\vec{F} = m\vec{a}"

Projection on OX:


"mg\sin(\alpha)+F = ma\\
F = m(a-g\sin(\alpha))\\
a) \alpha= 30^\circ\\
F=90\cdot(1-9.81\cdot 0.5) = -351.45 N\\
b) \alpha = 60^\circ\\
F=90\cdot(1-9.81\cdot \frac{\sqrt{3}}{2}) = -674.61 N"